# ST-Think: How Multimodal Large Language Models Reason About 4D Worlds from Ego-Centric Videos

### v1

@@ Abstract
Humans excel at spatio-temporal reasoning, effortlessly interpreting dynamic visual events from an egocentric viewpoint. However, whether multimodal large language models (MLLMs) can similarly comprehend the 4D world remains uncertain. This paper explores multimodal spatio-temporal reasoning from an egocentric perspective, aiming to equip MLLMs with human-like reasoning capabilities. To support this objective, we introduce Ego-ST Bench, a novel benchmark containing over 5,000 question-answer pairs across four categories, systematically evaluating spatial, temporal, and integrated spatio-temporal reasoning. Additionally, we propose ST-R1 Video model, a video-based reasoning model that incorporates reverse thinking into its reinforcement learning process, significantly enhancing performance. We combine long-chain-of-thought (long-CoT) supervised fine-tuning with Group Relative Policy Optimization (GRPO) reinforcement learning, achieving notable improvements with limited high-quality data. Ego-ST Bench and ST-R1 provide valuable insights and resources for advancing video-based spatio-temporal reasoning research.
Figure 1: Top: The Ego-ST bench is characterized by Egocentric Spatio-Temporal reasoning questions, and two forward spatio-temporal reasoning tasks are chosen here as examples: Route Description and Direction change selection. Bottom left: Shows the performance of each model in a spatio-temporal reasoning task. Bottom right: Shows the results of the 4 training methods compared to baseline. In the radar chart, F stands for forward, and R stands for reverse.
@@ 1 Introduction
Multimodal large language models with first-person spatio-temporal reasoning mimic human perception by seamlessly integrating visual, auditory, and textual data. This advanced capability enables a deeper understanding of dynamic environments, thereby supporting highly accurate predictions and context-aware decisions in fields such as autonomous driving, robotics, and augmented reality. Their human-like reasoning fosters more reliable, adaptable, and robust AI systems capable of handling complex tasks. Nevertheless, whether current large language models truly possess such sophisticated spatiotemporal reasoning abilities remains an open question that warrants further investigation.
@@
Recently, this issue has garnered widespread attention. VSI-Bench[37] introduced a benchmark for spatial reasoning and found that spatial reasoning capabilities remain the primary bottleneck for MLLMs. However, it mainly focuses on spatial aspects such as object and room sizes, spatial relationships, and distances, while lacking evaluation data for spatiotemporal reasoning tasks, such as the trajectory of ego motion, environmental description along paths, and action prediction. Moreover, it does not assess the network’s reverse reasoning abilities. In terms of models, VADAR[23] proposed an agentic approach to decompose spatial reasoning tasks and perform reasoning incrementally, but this approach is limited to static spatial problems and still fails to address spatiotemporal reasoning. Meanwhile, SpacialCoT[16] introduced spatial coordinate bi-directional alignment and chain-of-thought spatial grounding to enhance spatial reasoning, yet it still faces challenges in handling spatiotemporal reasoning, such as issues with temporal alignment. Therefore, we require a new benchmark for spatiotemporal reasoning to support research in this area. Moreover, training a model capable of spatiotemporal reasoning remains an important and challenging task.
@@
In this paper, we first propose a novel Ego-ST bench designed to evaluate current multimodal large language models and support cutting-edge research on spatiotemporal reasoning from an egocentric perspective. This comprehensive dataset comprises over 5,000 meticulously annotated instances, including detailed question-answer pairs and various multiple-choice questions. Notably, it is the first dataset to annotate both forward and reverse reasoning, thereby facilitating rigorous evaluation and in-depth research on bidirectional reasoning mechanisms. In our Ego-ST bench, our aim is to comprehensively assess the advanced perceptual capabilities of MLLMs. Unlike previous benchmarks such as HourVideo [2] and VSI-Bench [37], our work places a stronger emphasis on accurately describing complete routes, reasoning about directional changes and landmarks, and eliciting more granular and precise model responses. Inspired by Lu’s work on LLMs [18] and the inherent challenge of retracing reasoning paths frequently encountered in everyday human experiences, we introduce the novel concept of Reverse Thinking for the first time in the domain of multimodal large-scale model ego-video analysis. This innovative approach encourages MLLMs to adopt a human-centric perspective when interpreting ego-video content, enabling us to rigorously evaluate their capabilities in both spatiotemporal recall and reverse reasoning. Moreover, our benchmark sets a new standard that drives further research and paves the way for significant advancements in this emerging field.
@@
In addition, we propose a new spatiotemporal reasoning model, ST-R1. Building on insights from reverse thinking, the chain-of-thought (CoT) concept [31], and reinforcement learning approaches exemplified by Deepseek R1 [10], we have developed novel strategies for training a robust spatiotemporal reasoning model. Our core idea is to employ reverse thinking as the thought process to enhance spatiotemporal reasoning capabilities. To the best of our knowledge, we are the first to adopt reverse spatiotemporal thinking to boost the spatiotemporal reasoning ability of multimodal models. As shown in Figure 5, our ST-R1 Video model restructures the tasks of forward and reverse route descriptions by using the latter as a reasoning process. Specifically, we construct spatiotemporal CoT data that enables the model to learn the reasoning style and logic through Supervised Fine-Tuning (SFT) while retaining key generic capabilities. Subsequently, we employ Group Relative Policy Optimization (GRPO) as a second stage to further enhance the model’s reasoning performance.
@@
On the Ego-ST Bench, we found that existing models still exhibit considerable room for improvement in spatiotemporal reasoning and question answering. Surprisingly, there is no significant performance gap between open-source and closed-source models in this domain; in fact, some open-source models supporting longer context windows even achieve superior performance. This suggests that there is substantial research potential for enhancing the spatiotemporal reasoning capabilities of these networks. By post-training with a small amount of high-quality long CoT data, we developed the ST-R1 model, which significantly outperforms traditional SFT methods. This outcome underscores the effectiveness of our proposed multi-stage post-training strategy.
@@
  • We propose the novel concept of Reverse Thinking in the context of ego-video analysis for multimodal models. This approach mimics human reasoning by retracing the reasoning path, providing a human-centric perspective that enhances spatiotemporal recall and reverse reasoning capabilities—an innovation not previously explored in this domain.
  • We develop a new spatiotemporal reasoning model, ST-R1, which leverages insights from reverse thinking, the CoT concept, and reinforcement learning strategies. Our multi-stage training process, starting with SFT using spatiotemporal CoT data and followed by GRPO, significantly improves reasoning performance compared to traditional methods.
@@ 2 Related Work
Egocentric Video Understanding. Egocentric video understanding is a critical subfield of video analysis that provides a promising avenue for investigating whether multi-modal large language models (MLLMs) can emulate human-like reasoning. Moreover, it offers valuable insights for research in embodied intelligence and robotics. Compared to traditional video understanding tasks—such as those involving MSVD-QA [35], MSRVTT-QA [35], ActivityNet-QA [38], and LongVideoBench [32] egocentric video understanding is inherently more complex, demanding greater time and computational resources. In response to these challenges, numerous scholars have introduced several pivotal datasets, including Ego4D [9], EPIC-KITCHEN [7], HD-EPIC [26], and HOI4D [17]. Building upon these resources, researchers have recently begun to explore the capacity of MLLMs in the context of egocentric video, leading to the development of several benchmarks such as HourVideo [3], EgoPlan-Bench2 [28], X-LeBench [39], Egoschema[22], AMEGO[8] and VMB[11]. However, none of these benchmarks specifically target the evaluation of forward and reverse higher-level spatio-temporal reasoning capabilities in MLLMs. To address this gap, we propose the Ego-ST bench, which offers a detailed evaluation and analysis framework designed to assess these advanced reasoning skills.
@@
Multimodal Spatial Benchmarks. In previous studies, numerous scholars have evaluated the spatial capabilities of MLLMs. For instance, SpatialVLM [4] and SpatialRGPT [6] primarily address 2D image-level spatial understanding. However, spatio-temporal comprehension is critical in the video domain. Although benchmarks such as HourVideo [2], OVO-Bench [15], and X-LeBench [39] have assessed spatio-temporal understanding, an in-depth evaluation of multimodal spatio-temporal reasoning remains lacking. To address this gap, we propose the Ego-ST bench, which specifically focuses on evaluating higher-level spatio-temporal reasoning abilities in MLLMs.
@@
Mutimodal Reasoning. As large language models (LLMs) such as OpenAI-o1 [24] and DeepSeek-R1 [10] continue to gain prominence, a growing body of research has explored their application in multimodal domains. Notable examples include multimodal frameworks like LLava-o1 [36] and Open-R1-Multimodal [14], which seek to bridge the gap between textual, visual, and other data modalities. However, we contend that the challenges in multimodal domains do not currently require the same chain-of-thought reasoning techniques that have proven effective for mathematics, coding, and logic. Rather, the critical need lies in addressing advanced spatio-temporal reasoning problems that traditional methods struggle to resolve. In response to this gap, we introduce the ST-R1 model, which is specifically designed to excel in complex spatio-temporal reasoning tasks. By leveraging recent advances in multimodal reasoning and focusing on the intricate interplay between spatial and temporal features, the ST-R1 model aims to push the boundaries of multimodal reasoning and establish a new benchmark for tackling these sophisticated challenges.
@@ 3.1 Overview
We present the Ego-ST bench, which evaluate higher-level spatio-temporal reasoning abilities in egocentric video-based MLLM. The Ego-ST bench consists of over 5000 Question and Answer pairs from 789 real video clips. In total, these videos are divided into two types of scenarios, which include indoor scenarios and ever-changing outdoor complex scenarios. And there are five video sources, including video clips from SUN3D [34], HUJI [27], DoMSEV [29], Project Aria[19, 19], and self-collection video data. The reuse of these existing ego-video datasets provides accurate and difficult annotations. The Ego-ST bench consists of eight tasks of four types: Route Description, Direction Change Description and Action Change Description. Each of these types of tasks contains problem settings for forward and reverse reasoning, fully assessing the capabilities of MLLM in spatio-temporal reasoning. We believe that only a model that truly has the ability to reason inversely about forward video information can be called capable of spatio-temporal reasoning. For an overview of the Ego-ST bench task with statistical information about the dataset, see Figure 3. For related statistics, please refer to Figure 4.
@@ 3.2 Benchmark Construction
Data Collection and Unification. We conducted detailed research and screening of past egocentric video datasets, and finally, we selected 6 egocentric video datasets, including SUN3D [34], HUJI [27], DoMSEV [29], Aria Everyday Activities [19], Aria Digital Twin [25], and Nymeria [20]. In addition to this, we also shot a portion of the video to make up for the lack of richness of the picture in the locations where the original video data was concentrated. In these data we collated the describable parts by manually filtering and cropping the clips to make sure that the content of each video clip is labellable and has spatio-temporal reasoning implications. In the end, we collated and filtered a total of 789 video clips with videos containing indoor single scenes, multi-scenes and outdoor open complex scenes.
Question-Answer Generation. We spent a total of over 150 hours in manual filtering and annotation time, from video data filtering to QA pair creation. We carefully designed and refined the labelling templates for the positive and negative questions for the videos of four different scenarios (2 outdoor scenes and 2 indoor scenes). Among them, as it involves reverse route annotation, it requires the annotator to have a very high concentration and reverse spatial-temporal reasoning ability.
Based on the positive and negative route annotation information of the video data of the four sections, we elaborated a set of automated multiple-choice question generation and calibration pipeline, as shown in Figure 2. Based on this pipeline, we use the gpt-4o to generate three types of multiple-choice questions (direction description, landmark description, and action description) from the annotated data. Details of the api prompt we used can be found in Appendix.
@@
4 Spatial-Temporal R1 Model
The Spatial-Temporal R1 (ST-R1) model is a framework designed to understand and reason about 4D spatio-temporal information in videos. The goal of ST-R1 is to interpret complex events in videos by tracking objects, actions, and their interactions across space and time. By modeling both spatial configurations and temporal dynamics, ST-R1 can comprehend scenarios such as physical events, human activities, and scene changes, enabling it to answer questions or make inferences about video content. Formally, given a video (a sequence of frames) and a query, ST-R1 produces a reasoned answer by analyzing visual cues through time. To achieve this high-level understanding, as shown in Figure 5 ST-R1 is trained in two stages: first with supervised learning to teach chain-of-thought reasoning and second with reinforcement learning to refine its policy for video question-answering.
@@ 4.1 Chain-of-Thought Supervised Fine-Tuning
where $\theta$ are the model parameters. This loss expands over the tokens of the chain-of-thought and answer, ensuring the model learns to think aloud and arrive at the correct answer step by step. We use the forward problem as the thinking process for the backward problem and the backward problem as the thinking process for the forward problem.
Integrating forward and backward reasoning has been shown to improve overall reasoning accuracy[12], as the model learns not only to derive answers but also to validate them. After this CoT fine-tuning stage, ST-R1 has a strong initial ability to produce reasoned answers for video-based questions, but there is room to optimize its performance and alignment further. This motivates our second training stage using reinforcement learning.
Figure 5: Spatial Temporal Reasoning Model. Our model is trained in two stages: (1) Create Chain of Think (CoT) data for supervised fine-tuning (SFT). (2) Enhancing the model using the rule-based reinforcement learning GRPO algorithm.

### v2

@@ Abstract
Humans excel at spatial-temporal reasoning, effortlessly interpreting dynamic visual events from an egocentric viewpoint. However, whether multimodal large language models (MLLMs) can similarly understand the 4D world remains uncertain. This paper explores multimodal spatial-temporal reasoning from an egocentric perspective, aiming to equip MLLMs with human-like reasoning capabilities. To support this objective, we introduce Ego-ST Bench, a novel benchmark containing over 5,000 question-answer pairs across four categories, systematically evaluating spatial, temporal, and integrated spatial-temporal reasoning. Additionally, we propose ST-R1 training paradigm, a video-based reasoning model that incorporates reverse thinking into its reinforcement learning process, significantly enhancing performance. We combine long-chain-of-thought (long-CoT) supervised fine-tuning with Group Relative Policy Optimization (GRPO) reinforcement learning, achieving notable improvements with limited high-quality data. Ego-ST Bench and ST-R1 provide valuable insights and resources for advancing video-based spatial-temporal reasoning research.
Figure 1: Top: The Ego-ST bench is characterized by Egocentric Spatio-Temporal reasoning questions, and two spatial-temporal reasoning tasks are chosen here as examples: Route Description and Direction change selection. Bottom left: Shows the performance of each model in a spatial-temporal reasoning task. Bottom right: Shows the results of the 4 training methods compared to baseline. (In the radar chart, F stands for forward and R stands for reverse.)
@@ 1 Introduction
Multimodal large language models with first-person spatial-temporal reasoning mimic human perception by seamlessly integrating visual, auditory, and textual data. This advanced capability enables a deeper understanding of dynamic environments, thereby supporting highly accurate predictions and context-aware decisions in fields such as autonomous driving, robotics, and augmented reality. Their human-like reasoning fosters more reliable, adaptable, and robust AI systems capable of handling complex tasks. Nevertheless, whether current large language models truly possess such sophisticated spatialtemporal reasoning abilities remains an open question that warrants further investigation.
@@
Recently, this issue has garnered widespread attention. VSI-Bench[34] introduced a benchmark for spatial reasoning and found that spatial reasoning capabilities remain the primary bottleneck for MLLMs. However, it mainly focuses on spatial aspects such as object and room sizes, spatial relationships, and distances, while lacking evaluation data for spatialtemporal reasoning tasks, such as the trajectory of ego motion, environmental description along paths, and action prediction. Moreover, it does not evaluate the network’s reverse reasoning abilities. In terms of models, VADAR[20] proposed an agentic approach to decompose spatial reasoning tasks and perform reasoning incrementally, but this approach is limited to static spatial problems and still fails to address spatialtemporal reasoning. Meanwhile, SpacialCoT[14] introduced spatial coordinate bi-directional alignment and chain-of-thought spatial grounding to enhance spatial reasoning, yet it still faces challenges in handling spatialtemporal reasoning, such as issues with temporal alignment. Therefore, we require a new benchmark for spatialtemporal reasoning to support research in this area. Moreover, training a model capable of spatialtemporal reasoning remains an important and challenging task.
@@
In this paper, we first propose a novel Ego-ST bench designed to evaluate current multimodal large language models and support cutting-edge research on spatialtemporal reasoning from an egocentric perspective. This comprehensive dataset comprises over 5,000 meticulously annotated instances, including detailed question-answer pairs and various multiple-choice questions. Notably, it is the first dataset to annotate both forward and reverse reasoning, thereby facilitating rigorous evaluation and in-depth research on bidirectional reasoning mechanisms. In our Ego-ST bench, our aim is to comprehensively evaluate the advanced perceptual capabilities of MLLMs. Unlike previous benchmarks such as HourVideo [2] and VSI-Bench [34], our work places a stronger emphasis on accurately describing complete routes, reasoning about directional changes and landmarks, and eliciting more granular and precise model responses. Inspired by Lu’s work on LLMs [16] and the inherent challenge of retracing reasoning paths frequently encountered in everyday human experiences, we introduce the novel concept of Reverse Thinking for the first time in the domain of multimodal large-scale model ego-video analysis. This innovative approach encourages MLLMs to adopt a human-centric perspective when interpreting ego-video content, enabling us to rigorously evaluate their capabilities in both spatialtemporal recall and reverse reasoning. Moreover, our benchmark sets a new standard that drives further research and paves the way for significant advancements in this emerging field.
@@
In addition, we propose a new spatial-temporal reasoning training paradigm, ST-R1. Building on insights from reverse thinking, the chain-of-thought (CoT) concept [28], and reinforcement learning approaches exemplified by Deepseek R1 [9], we have developed novel strategies for training a robust spatialtemporal reasoning model. Our core idea is to employ reverse thinking as the thought process to enhance spatialtemporal reasoning capabilities. To the best of our knowledge, we are the first to adopt reverse spatialtemporal thinking to boost the spatialtemporal reasoning ability of multimodal models. As shown in Figure 5, our ST-R1 Video model restructures the tasks of forward and reverse route descriptions by using the latter as a reasoning process. Specifically, we construct spatialtemporal CoT data that enables the model to learn the reasoning style and logic through Supervised Fine-Tuning (SFT) while retaining key generic capabilities. Subsequently, we employ Group Relative Policy Optimization (GRPO) as a second stage to further enhance the model’s reasoning performance.
@@
On the Ego-ST Bench, we found that existing models still exhibit considerable room for improvement in spatialtemporal reasoning and question answering. Surprisingly, there is no significant performance gap between open-source and closed-source models in this domain; in fact, some open-source models supporting longer context windows even achieve superior performance. This suggests that there is substantial research potential for enhancing the spatialtemporal reasoning capabilities of these networks. By post-training with a small amount of high-quality long CoT data, we developed the ST-R1 model, which significantly outperforms traditional SFT methods. This outcome underscores the effectiveness of our proposed multi-stage post-training strategy.
@@
  • We propose the novel concept of Reverse Thinking in the context of ego-video analysis for multimodal models. This approach mimics human reasoning by retracing the reasoning path, providing a human-centric perspective that enhances spatialtemporal recall and reverse reasoning capabilities—an innovation not previously explored in this domain.
  • We develop a new spatialtemporal reasoning training paradigm, ST-R1, which leverages insights from reverse thinking, the CoT concept, and reinforcement learning strategies. Our multi-stage training process, starting with SFT using spatialtemporal CoT data and followed by GRPO, significantly improves reasoning performance compared to traditional methods.
@@ 2 Related Work
Egocentric Video Understanding. Egocentric video understanding is a critical subfield of video analysis that provides a promising avenue for investigating whether multi-modal large language models (MLLMs) can emulate human-like reasoning. Moreover, it offers valuable insights for research in embodied intelligence and robotics. Compared to traditional video understanding tasks—such as those involving MSVD-QA [32], MSRVTT-QA [32], ActivityNet-QA [35], and LongVideoBench [29] egocentric video understanding is inherently more complex, demanding greater time and computational resources. In response to these challenges, numerous scholars have introduced several pivotal datasets, including Ego4D [8], EPIC-KITCHEN [7], HD-EPIC [23], and HOI4D [15]. Building upon these resources, researchers have recently begun to explore the capacity of MLLMs in the context of egocentric video, leading to the development of several benchmarks such as HourVideo [3], EgoPlan-Bench2 [25], and X-LeBench [36]. However, none of these benchmarks specifically target the evaluation of higher-level spatial-temporal reasoning capabilities in MLLMs. To address this gap, we propose the Ego-ST bench, which offers a detailed evaluation and analysis framework designed to evaluate these advanced reasoning skills.
@@
Multimodal Spatial Benchmarks. In previous studies, numerous scholars have evaluated the spatial capabilities of MLLMs. For instance, SpatialVLM [4] and SpatialRGPT [6] primarily address 2D image-level spatial understanding. However, spatial-temporal comprehension is critical in the video domain. Although benchmarks such as HourVideo [2], OVO-Bench [13], and X-LeBench [36] have assessed spatial-temporal understanding, an in-depth evaluation of multimodal spatial-temporal reasoning remains lacking. To address this gap, we propose the Ego-ST bench, which specifically focuses on evaluating higher-level spatial-temporal reasoning abilities in MLLMs.
@@
Mutimodal Reasoning. As large language models (LLMs) such as OpenAI’s O1 [21] and DeepSeek-R1 [9] continue to gain prominence, a growing body of research has explored their application in multimodal domains. Notable examples include multimodal frameworks like LLava-O1 [33] and Open-R1-Multimodal [12], which seek to bridge the gap between textual, visual, and other data modalities. However, we contend that the challenges in multimodal domains do not currently require the same chain-of-thought reasoning techniques that have proven effective for mathematics, coding, and logic. Rather, the critical need lies in addressing advanced spatial-temporal reasoning problems that traditional methods struggle to resolve. In response to this gap, we introduce the ST-R1 model, which is specifically designed to excel in complex spatial-temporal reasoning tasks. By leveraging recent advances in multimodal reasoning and focusing on the intricate interplay between spatial and temporal features, the ST-R1 model aims to push the boundaries of multimodal reasoning and establish a new benchmark for tackling these sophisticated challenges.
@@ 3.1 Overview
We present the Ego-ST bench, which evaluate higher-level spatial-temporal reasoning abilities in egocentric video-based MLLM. The Ego-ST bench consists of over 5000 Question and Answer pairs from 789 real video clips. In total, these videos are divided into two types of scenarios, which include indoor scenarios and ever-changing outdoor complex scenarios. And there are five video sources, including video clips from SUN3D [31], HUJI [24], DoMSEV [26], Aria Project and self collection video data. The reuse of these existing ego-video datasets provides accurate and difficult annotations. The Ego-ST bench consists of eight tasks of four types: Route Description, Direction Change Description and Action Change Description. Each of these types of tasks contains problem settings for forward and reverse reasoning, fully assessing the capabilities of MLLM in spatial-temporal reasoning. We believe that only a model that truly has the ability to reason inversely about forward video information can be called capable of spatial-temporal reasoning. For an overview of the Ego-ST bench task samples, see Figure 3. For related statistics, please refer to Figure 4.
@@ 3.2 Benchmark Construction
Data Collection and Unification. We conducted a detailed research and screening of past egocentric video datasets, and finally we selected 6 egocentric video datasets, including: SUN3D [31], HUJI [24] , DoMSEV [26], Aria Everyday Activities [17], Aria Digital Twin [22] and Nymeria [18]. In addition to this, we also collect a portion of the video to make up for the lack of richness of the picture in the locations where the original video data was concentrated. In these data we collated the describable parts by manually filtering and cropping the clips to make sure that the content of each video clip is labellable and has spatial-temporal reasoning implications. In the end, we collated and filtered a total of 789 video clips with videos containing indoor single scenes, multi-scenes and outdoor open complex scenes.
Question-Answer Generation. We spent a total of over 150 hours (each person) in manual filtering and annotation time, from video data filtering to QA pair creation. We carefully designed and refined the labelling templates for the positive and negative questions for the videos of four different scenarios (2 outdoor scenes and 2 indoor scenes). Among them, as it involves reverse route annotation, it requires the annotator to have a very high concentration and reverse spatial-temporal reasoning ability.
Based on the forward and reverse route annotation information of the video data of the four sections, we elaborated a set of automated multiple-choice question generation and calibration pipeline, as shown in Figure 2. Based on this pipeline, we use the gpt-4o to generate three types of multiple-choice questions (direction description, landmark description, and action description) from the annotated data.
@@
4 Spatial-Temporal R1
Next we will introduce a new paradigm for training in spatial-temporal tasks. The Spatial-Temporal R1 (ST-R1) model is a framework designed to understand and reason about 4D spatial-temporal information in videos. The goal of ST-R1 is to interpret complex events in videos by tracking objects, actions, and their interactions across space and time. By modeling both spatial configurations and temporal dynamics, ST-R1 can comprehend scenarios such as physical events, human activities, and scene changes, enabling it to answer questions or make inferences about video content. Formally, given a video (a sequence of frames) and a query, ST-R1 produces a reasoned answer by analyzing visual cues through time. To achieve this high-level understanding, as shown in Figure 5 ST-R1 is trained in two stages: first with supervised learning to teach chain-of-thought reasoning and second with reinforcement learning to refine its policy for video question-answering.
@@ 4.1 Chain-of-Thought Supervised Fine-Tuning
where $\theta$ are the model parameters. This loss expands over the tokens of the chain-of-thought and answer, ensuring the model learns to think aloud and arrive at the correct answer step by step. We use the forward problem as the thinking process for the reverse problem and the reverse problem as the thinking process for the forward problem.
Integrating forward and reverse reasoning has been shown to improve overall reasoning accuracy[10], as the model learns not only to derive answers but also to validate them. After this CoT fine-tuning stage, ST-R1 has a strong initial ability to produce reasoned answers for video-based questions, but there is room to optimize its performance and alignment further. This motivates our second training stage using reinforcement learning.
Figure 5: Spatial Temporal Reasoning Model. Our model is trained in two stages: (1) Create Chain of Thought (CoT) data for supervised fine-tuning (SFT). (2) Enhancing the model using the rule-based reinforcement learning GRPO algorithm.